A gust of wind blows an apple from a tree.
As the apple falls, the force of gravity on the
apple is 9.22 N downward, and the force of
the wind on the apple is 1.94 N to the right.
What is the magnitude of the net external
force on the apple?
Answer in units of N
1
Expert's answer
2019-10-14T09:31:57-0400


The net external force on the apple


"F=\sqrt{F_1^2+F_2^2}=\sqrt{9.22^2+1.94^2}=9.42\:\rm N"